IU switch migration led to network issues IU <-> UC and IU <-> CERN over the last week. Seems to be fixed now. Taking the time to rebuild all IU worker nodes as EL9.
Large batch of jobs were OOMing workers at UC yesterday (Mar 5). Identified the task and submitted a report to DPD. Rod Walker halted the jobs and we're recovering workers at UC now. Taking the opportunity to rebuild machines that went down as EL9.
Cooling event in a data center at UC brought down a few services that the AF and rebuilding nodes use. Recovered quickly, and now back to full functionality in that regard.
Rebuilding workers and storage at UC as EL9 to meet the June 30th EoL.
UIUC data center migration pushed back to a currently unknown date. We are inquiring about the status of EL9 for the nodes.
DC24 pushed our older storage to almost max, requested some deletions of DC24 data to make sure production is not effected. Made adjustments to some of our older storage, as well, to handle the amount of data flow. Hoping to retire said storage with our next purchase.